The data below were collected for the following reaction at 500 ∘C:
CH3CN(g)→CH3NC(g)
Time (h) [CH3CN] (M)
0.0 1.000
5.0 0.794
10.0 0.631
15.0 0.501
20.0 0.398
25.0 0.316

Part A: Determine the order of the reaction at this temperature.
A) zero order
B) first order
C) second order
D) third order

Part B: What is the half-life for this reaction (at the initial concentration)?
t1/2 = ?hr

Part C: How long will it take for 82% of the CH3CN to convert to CH3NC?
t= ?hr
1
Expert's answer
2020-06-11T09:22:15-0400

Part A. B) first order, lg(C0/C) = 0.02t

Part B. t1/2 = 0.693 / 0.02 = 34 min = 0.58 hr

Part C. t = lg(C0 / C) / 0.02 = lg(1 / (1 - 0.82)) / 0.02 = 37.2 min = 0.62 hr.